### v1

@@ Abstract
Temporal knowledge graph reasoning (TKGR) is increasingly gaining attention for its ability to extrapolate new events from historical data, thereby enriching the inherently incomplete temporal knowledge graphs. Existing graph-based representation learning frameworks have made significant strides in developing evolving representations for both entities and relational embeddings. Despite these achievements, there’s a notable tendency in these models to inadvertently learn biased data representations and mine spurious correlations, consequently failing to discern the causal relationships between events. This often leads to incorrect predictions based on these false correlations. To address this, we propose an innovative causal enhanced graph representation learning framework for TKGR (named CEGRL-TKGR). This framework introduces causal structures in graph-based representation learning to unveil the essential causal relationships between events, ultimately enhancing task performance. Specifically, we first disentangle the evolutionary representations of entities and relations in a temporal graph sequence into two distinct components, namely causal representations and confounding representations. Then, drawing on causal intervention theory, we advocate the utilization of causal representations for predictions, aiming to mitigate the effects of erroneous correlations caused by confounding features, thus achieving more robust and accurate predictions. Finally, extensive experimental results on six benchmark datasets demonstrate the superior performance of our model in the link prediction task.
@@ 1 Introduction
Knowledge graphs (KGs) have gained significant promise in natural language processing or knowledge engineering perception tasks [5]. They model real-world fact information by using multi-relationship graph structures. However, factual knowledge in reality is constantly evolving, resulting in the form of event knowledge. This has led to the development and application of temporal knowledge graphs (TKGs). TKGs encode the relationship information of entities and events and their timing. This enables TKGs to capture the dynamics of entities and their relationships over time [11]. Thus, analyzing TKGs provides a comprehensive understanding of the evolving events, based on which various time-dependent applications have been developed, including time-sensitive semantic search [1], policy making [8], stock forecasting [9], and more [5].
@@
The reliability of applications depends on accurate predicting, which highly relies on data integrality. However, existing TKGs are inevitably incomplete due to the partial observation of real-world [22]. To address this limitation and enhance the capability of TKGs, temporal knowledge graph reasoning (TKGR) models are proposed. TKGR models extrapolate new facts and relationships of TKGs according to their historical temporal information. Existing TKGR models explored different strategies to achieve satisfactory results on TKG extrapolation tasks. GHNN [13] and GHT [28] model historical facts as point-in-time processes. RE-NET [15] and RE-GCN [21] introduce graph neural networks (GNN) into sequence models to capture structural and temporal dependencies between entities. TLogic [23] and TITer [29] design interpretable models based on logical rules and reinforcement learning, respectively.
Figure 1: An illustrative example to explain that confounding factors such as trivial patterns might mislead the link prediction.
Despite the achievements of previous studies, they have overlooked the reality that there are numerous confounding factors in TKGs, such as shallow patterns and noisy links. However, these confounding factors commonly misguide the reasoning process of TKGs, resulting in the acquisition of incorrect dependencies and the generation of non-causal predictions [27]. An illustrative example of this phenomenon present in policy making is depicted in Fig. 1. Typically, a treaty between two countries is formed depending on months and years of negotiations. Nevertheless, there are instances where a state visit may occur between the last negotiation session and the finalization of the treaty. In such a scenario, the state visit becomes a confounding factor influencing the treaty’s finalization, since the consequences of a state visit are various and the treaty could only be finalized according to multiple rounds of negotiations. Nevertheless, a TKGR model would misinterpret the state visit as an indication of treaty finalization, based on limited observations of their simultaneous occurrences.
@@
To address the aforementioned issue, we advocate for the integration of causal theory into TKGR to guide learning of the essential causal relationships between events and mitigate the impact of confounding factors on TKGR. Specifically, we first construct a structural causal model [34] to comprehensively analyze and model TKGR tasks from a causal perspective. Then, based on the causal model, we propose a new framework, namely causal enhanced graph representation learning (CEGRL-TKGR), to disentangle confounding factors from the essential causal factors in TKGs. The framework utilizes graph neural networks and causal intervention techniques to learn the dynamic To the best of our knowledge, this is the first study to incorporate causal intervention in a graph representation learning framework for learning the evolutionary representations of entities and relations for TKG extrapolation. To conclude, our contributions are 4-folds:
  • To the best of our knowledge, this is the first study to incorporate causal intervention in a graph representation learning framework for learning the evolutionary representations of entities and relations in TKGs.
@@
  • We propose a novel causal enhanced graph representation learning framework for TGKR, called CEGRL-TKGR, to uncover the essential causal relationships between events and mitigate the impact of confounding factors.
@@
  • The proposed CEGRL-TKGR model disentangles the evolutionary representations of entities and relations into causal and confounding representations. Then, it applies causal interventions to perform backdoor adjustments of representations, prioritizing predicted causal features while minimizing the impact of spurious correlations introduced by confounding features.
@@
  • Comprehensive experimental results demonstrate the CEGRL-TKGR model outperforms state-of-the-art baselines on six real-world datasets in the link prediction task. Further, comprehensive studies confirm the contribution of the introduced causal structures and interventions [^1] Our code will be publicly available once the work is accepted..
@@ 2 Related Work
In this section, we briefly review the research advance of several works that are closely related to our work.
@@ 2.1 Temporal Knowledge Graph Reasoning
TKG inference in extrapolation settings focuses on predicting new facts about the future based on historical events. Specifically, CyGNet [37] uses a copy-generating mechanism to capture the global repetition rate of facts. GHNN [13] and GHT [28] construct a temporal point process (TPP) to capture the temporal dynamics of successive events, predicting future facts by estimating the conditional probability of the TPP. In recent years, with the successful application of GNN in many dynamic scenarios [35], they have also been introduced into structural-semantic dependency models in TKGR. RE-NET [15] used a neighborhood aggregator and cyclic event encoder to model historical facts as subgraph sequences. RE-GCN [21] uses RGCN [26] to learn evolutionary representations of entities and relationships at each timestamp. CEN [19] uses length-aware convolutional neural networks (CNNS) to process evolutionary patterns of different lengths. There are also some studies to solve the TKG inference problem through path search. For example, TLogic [23] completes link prediction tasks based on temporal logic rules learned from temporal knowledge graphs. TITer [29] proposes a TKG prediction model based on reinforcement learning, which uses time-shaped rewards based on Dirichlet distribution to guide model training. All of the methods discussed above have limitations in modeling entity and relationship representations, in particular ignoring cause-and-effect relationships between different entities, which we believe is key to making correct predictions.
@@ 2.2 Causal Representation Learning
In graph causal representation learning, researchers have explored various methods to improve the explanatory power and generalization performance of GNNs. By applying the principles of causal reasoning to graph-structured data, the researchers sought to address the challenges GNNs face when dealing with complex systems such as social networks, molecular maps, and syntax trees of program code. DIR [32] is proposed to reveal the intrinsic interpretability of GNNs by discovering invariant reasons, which involves splitting input graphs into causal and non-causal fruit graphs and training the two classifiers through invariant risk loss functions. GOOD [6] improves the cross-domain generalization of graphs by distinguishing invariant subgraphs from other parts of graphs that are susceptible to domain transfer. CAL [27] introduces de-confounding training to distinguish the key and secondary parts of the graph and eliminate the confounding effect of the secondary parts on model prediction. CFLP [36] points out that the causal relationship between graph structure and link presence is often ignored, and proposed to generate counterfactual links to enhance training data and reduce reliance on false associations. Zevcevic et al. [34] theoretically analyzed the relationship between GNNs and structural causal models (SCMs) and designed a new class of neuro-causal models. However, none of the work has been done to combine causal learning with temporal knowledge graph reasoning.
@@ 3.1 Notations and Task Formulation
A TKG $\mathcal{G}$ can be formalized as a sequence of graph slices $\left\{\mathcal{G}_{1},\mathcal{G}_{2},\ldots.,\mathcal{G}_{T}\right\}$, where $\mathcal{G}_{t}=\left\{\left(e_{s},r,e_{o},t\right)\in\mathcal{G}\right\}$ denotes a graph slice that consists of facts that occurred at the timestamp $t$. Here, $e_{s}$ and $e_{o}$ represent the subject and object entities, respectively, and $r$ denotes the predicate as a relation type. Besides, $\mathbf{e}_{s}$, $\mathbf{r}$, $\mathbf{e}_{o}$ written in bold represent their embeddings. Our goal is to predict the future timestamp of the query missing entity. The task can be defined as follows: Given a query $\left(e_{s},r,?,t\right)$, the model needs to predict the missing entity.
3.2 A Causal Perspective on the TKGR Task
To deepen the comprehension of the GNN-based TKG inference mechanism, we abstract the GNN-based inference process through a causal figure, as shown in Fig. 2, encompassing five distinct variables. The connectivity from one variable to another epitomizes the causal relationship, delineated as the cause $\rightarrow$ effect. The variables are described as follows:
Figure 2: The GNN-based causal graph on TKG inference process.
  • Graph data $\mathcal{G}_{t}$: The knowledge graph data at each timestamp t, manifests as a directed relationship figure.
@@ 3.2 A Causal Perspective on the TKGR Task
  • Prediction $Y$: Denoted as the link prediction, this aspect transitions through the decoder, rendering the ultimate inference based on the preceding representation.
The causal embedding encapsulates the causal features $C$, authentically mirroring the implicit knowledge inherent in the knowledge graph data $\mathcal{G}_{t}$. Conversely, $N$ symbolizes the confounding features, which might be spawned by data biases, data noise, or superficial patterns within graph-based learning methodologies. These confounding features forge a backdoor pathway between $C$ and $Y$, fostering spurious correlations that don’t contribute to accurate inference. Functionally, the structural operation denoted by $C$rightarrow $R$leftarrow $N$ portrays a GNN, wherein both the causal features $C$ and the confounding features $N$, as discerned by the target entity from the graph data, exert a direct impact on the output $R$ of the GNN. Subsequently, the output $R$ of GNN directly sways the model inference outcome, illustrated as $R$rightarrow $Y$.
In the graph-based TKGR paradigm, causal and confounding features are not decoupled for each entity or relationship embedding. Using causal graphs, we aspire to explicitly separate causal embeddings and confounding embeddings from entity or relational representations, and aim to mitigate the effects of confounding features by performing causal interventions. This endeavor not only clarifies the inference process but also endeavors to refine the accuracy and reliability of the GNN-based TKG inference mechanism.
3.3 Causal Intervention
Beyond fostering a novel comprehension of GNN-based TKGR, causal theory avails analytical instruments predicated on causal figures, such as causal intervention. Causal intervention facilitates a profound examination of the factors precipitating inference outcomes. As delineated by the causal figure, confounding feature $N$ and causal feature $C$ can be discerned from the knowledge graph data. These features are contemplated in the representation R of entities and relations, thereby establishing a backdoor pathway represented as $N$leftarrow $\mathcal{G}_{t}$rightarrow $C$rightarrow $R$rightarrow $Y$, with $N$ serving as the quick bridge between $C$ and $Y$.
@@ 3.3 Causal Intervention
$P(Y\mid do(C))=\sum_{n\in N}P(Y\mid C,n)P(n)$ (1)
The equation above illustrates that to gauge the causal influence of $C$ on $Y$, it’s requisite to take into account the inference outcomes of both causal and confounding features. This can be perceived as re-coupling the disentanglement feature embeddings, utilizing them for deductive reasoning at future timestamps. However, $C$ and $N$ are usually unobservable, and it is difficult to obtain them directly at the data level, which makes the calculation of the Eq. (1) very challenging. In the next section, we devote ourselves to discussing ways to overcome this problem.
4 The Proposed Approach
In this section, we detail CEGRL-TKGR, the proposed causal enhancement TKGR framework for learning representations of entities and relationships based on causal features and confounding features. CEGRL-TKGR model consists of three parts: the representation learning part that learns the structure dependence in each $\mathcal{G}_{t}$, the decoupling learning part that learns the entity and relation representations, and the decoder based on the time interval. The structure of the model framework is shown in Fig. 3.
Figure 3: The overall architecture of our proposed CEGRL-TKGR model.
@@ 4.1 Entity and Relation Evolution Representation
Within each $\mathcal{G}_{t}$, representation learning of entities and relationships involves the aggregation of multiple relationships, as well as information from multiple hop neighbors under a single timestamp. Between adjacent $\mathcal{G}_{t}$, we expect to accurately capture the order dependencies inherent in the subgraph with different timestamps. Drawing inspiration from the RE-GCN model [21], we employ the $\omega$-layer RGCN, which hinges on structure modeling and a recurrent mechanism to progressively update the representations of entities and relations. This approach allows for a more nuanced understanding and modeling of the dynamic interactions within the graph over time.
$\mathbf{e}_{o,t}^{l+1}$ $=\operatorname{RReLu}\left(\sum_{(e_{s},r,e_{o})\in\mathcal{G}_{t}}\frac{1}{d_{e_{o}}}\mathbf{W}_{1}^{l}\left(\Phi\left(\mathbf{e}_{s,t}^{l},\mathbf{r}_{t}\right)\right)+\mathbf{W}_{2}^{l}\mathbf{e}_{o,t}^{l}\right),$ (2)
$\mathbf{E}_{t}$ $=\operatorname{GRU}\left(\mathbf{E}_{t-1},\mathbf{E}_{t}^{\prime}\right).$ (3)
In the Eq. (2), we describe how the embedding $\mathbf{e}_{o,t}^{l+1}$ of entity $e_{o}$ at time step $t$ and layer $l+1$ is computed. We integrate the information of all entities and relations connected to entity $e_{o}$ in the graph $\mathcal{G}_{t}$. $\mathbf{W}_{1}^{l},\mathbf{W}_{2}^{l}$ is learnable weights and $\Phi$ has the option of addition or one-dimensional convolution. In the Eq. (3), we showcase how the entity embedding matrix $\mathbf{E}_{t}$ is updated via the GRU. Specifically, we take the entity embedding matrix $\mathbf{E}_{t-1}$ at the previous time step $t-1$ and the aggregated entity embedding matrix $\mathbf{E}_{t}^{\prime}$ as inputs to obtain the entity embedding matrix $\mathbf{E}_{t}$ at the current time step $t$.
@@
$\mathbf{R}_{t}$ $=\operatorname{GRU}\left(\mathbf{R}_{t-1},\mathbf{R}_{t}^{\prime}\right),$ (5)
where $\mathbf{r}_{t}^{\prime}$ is an aggregation of all entities connected to relation r via a mean pooling operation, and $\mathbf{R}_{t}^{\prime}$ is obtained by concatenating this result with the embeddings of all relations. Eventually, we update the relation embedding matrix $\mathbf{R}_{t}$ using a GRU.

### v2

@@ Abstract
Temporal knowledge graph reasoning (TKGR) is increasingly gaining attention for its ability to extrapolate new events from historical data, thereby enriching the inherently incomplete temporal knowledge graphs. Existing graph-based representation learning frameworks have made significant strides in developing evolving representations for both entities and relational embeddings. Despite these achievements, there’s a notable tendency in these models to inadvertently learn biased data representations and mine spurious correlations, consequently failing to discern the causal relationships between events. This often leads to incorrect predictions based on these false correlations. To address this, we propose an innovative Causal Enhanced Graph Representation Learning framework for TKGR (named CEGRL-TKGR). This framework introduces causal structures in graph-based representation learning to unveil the essential causal relationships between events, ultimately enhancing the performance of the TKGR task. Specifically, we first disentangle the evolutionary representations of entities and relations in a temporal knowledge graph sequence into two distinct components, namely causal representations and confounding representations. Then, drawing on causal intervention theory, we advocate the utilization of causal representations for predictions, aiming to mitigate the effects of erroneous correlations caused by confounding features, thus achieving more robust and accurate predictions. Finally, extensive experimental results on six benchmark datasets demonstrate the superior performance of our model in the link prediction task.
CEGRL-TKGR: A Causal Enhanced Graph Representation Learning Framework for Temporal Knowledge Graph Reasoning
Jinze Sun1, Yongpan Sheng1,2(✉), Lirong He3, Yongbin Qin4, Ming Liu5, Tao Jia1 1College of Computer and Information Science, Southwest University, Chongqing 400715, China 2School of Big Data & Software Engineering, Chongqing University, Chongqing 401331, China 3School of Information Science and Engineering, Chongqing Jiaotong University, Chongqing 400074, China 4College of Computer Science and Technology, The State Key Laboratory of Public Big Data, Guizhou University 5Deakin University, 3125, Victoria, Australia ssunjinze@outlook.com  {shengyp2011,ronghe1217}@gmail.com  ybqin@gzu.edu.cn  m.liu@deakin.edu.au  tjia@swu.edu.cn
@@ 1 Introduction
Knowledge graphs (KGs) have gained significant promise in natural language processing or knowledge engineering perception tasks Chen et al. (2022a). They model real-world factual knowledge using multi-relationship graph structures. However, factual knowledge in reality is constantly evolving, resulting in the form of event knowledge. This has led to the development and application of temporal knowledge graphs (TKGs). TKG encodes the relationship information of entities and events and their timing for capturing the dynamics of entities and their relationships over time Gastinger et al. (2022). Thus, analyzing the TKG provides a comprehensive understanding of the evolving events, based on which various time-dependent applications have been developed, including time-sensitive semantic search Barbosa et al. (2013), policy making Deng et al. (2020), stock forecasting Feng et al. (2019), and more Chen et al. (2022a).
@@
The reliability of applications depends on accurate predicting, which highly relies on data integrality. However, existing TKGs are inevitably incomplete due to the partial observation of real-world Liang et al. (2022). To address this limitation and enhance the representation capability of the TKG, temporal knowledge graph reasoning (TKGR) models are proposed and aim to extrapolate new facts and relationships in the TKG according to their historical temporal information. Existing models explore different strategies to achieve satisfactory results on the TKGR task. GHNN Han et al. (2020) and GHT Sun et al. (2022) model historical facts as point-in-time processes. TKGR-RHETNE Sun et al. (2023) jointly models the relevant historical event and temporal neighborhood event context of events in the TKG. RE-NET Jin et al. (2020) and RE-GCN Li et al. (2021) introduce graph neural networks (GNN) into sequence models to capture structural and temporal dependencies between entities. TKGR-GPRSCL Xiong et al. (2024) captures complex structure-aware information by encoding paths across entities and obtaining temporal correlations in the complex plane. TLogic Liu et al. (2022) and TITer Sun et al. (2021) design interpretable models based on logical rules and reinforcement learning, respectively. Despite the achievements of previous studies, they have overlooked the reality that there are numerous confounding factors in the TKG, such as shallow patterns and noisy links. However, these confounding factors commonly misguide the reasoning process in the TKG, resulting in the acquisition of incorrect dependencies and the generation of non-causal predictions Sui et al. (2022).
@@
To address the aforementioned issues, we advocate for the integration of causal theory into TKGR to guide learning of the essential causal relationships between events and mitigate the impact of confounding factors on the TKGR task. Specifically, we first construct a structural causal model Zečević et al. (2021) to comprehensively analyze and model the TKGR task from a causal perspective. Then, based on the causal model, we propose a new framework, namely Causal Enhanced Graph Representation Learning (CEGRL-TKGR), to disentangle confounding factors from the essential causal factors in the TKG. To the best of our knowledge, this is the first study to incorporate causal intervention in a graph representation learning framework for learning the evolutionary representations of entities and relations in the TKG. To conclude, our contributions in this paper are 3-folds:
@@
  • We propose a novel Causal Enhanced Graph Representation Learning framework for Temporal Knowledge Graph Reasoning, called CEGRL-TKGR, to uncover the essential causal relationships between events and mitigate the impact of confounding factors.
@@
  • The proposed CEGRL-TKGR framework disentangles the evolutionary representations of entities and relations into causal and confounding representations. Then, it applies causal interventions to perform backdoor adjustments of representations, prioritizing predicted causal features while minimizing the impact of spurious correlations introduced by confounding features.
@@
  • Comprehensive experimental results demonstrate that CEGRL-TKGR outperforms state-of-the-art baselines on six real-world datasets in the link prediction task. Further, comprehensive studies confirm the contribution of the introduced causal structures and interventions [^1] To illustrate the evaluation of our CEGRL-TKGR framework and facilitate further research on this topic, we have made the experimental details and source code of the framework publicly available at https://github.com/shengyp/CEGRL-TKGR..
@@ 2.1 Temporal Knowledge Graph Reasoning
TKGR in extrapolation settings focuses on predicting new facts about the future based on historical events. Specifically, CyGNet Zhu et al. (2021) uses a copy-generating mechanism to capture the global repetition rate of facts. GHNN Han et al. (2020) and GHT Sun et al. (2022) construct a temporal point process (TPP) to capture the temporal dynamics of successive events, predicting future facts by estimating the conditional probability of the TPP. In recent years, with the successful application of GNN in many dynamic scenarios Zhang et al. (2022), they have also been introduced into structural-semantic dependency models in TKGR. RE-NET Jin et al. (2020) used a neighborhood aggregator and cyclic event encoder to model historical facts as subgraph sequences. RE-GCN Li et al. (2021) uses RGCN Schlichtkrull et al. (2018) to learn evolutionary representations of entities and relationships at each timestamp. CEN Li et al. (2022a) uses length-aware convolutional neural networks (CNNS) to process evolutionary patterns of different lengths. There are also some studies to solve the TKGR problem through path search. For example, TLogic Liu et al. (2022) completes link prediction tasks based on temporal logic rules learned from temporal knowledge graphs. TITer Sun et al. (2021) proposes a TKG prediction model based on reinforcement learning, which uses time-shaped rewards based on Dirichlet distribution to guide model training. All of the methods discussed above have limitations in modeling entity and relationship representations, in particular ignoring cause-and-effect relationships between different entities, which we believe is key to making correct predictions.
@@ 2.2 Causal Representation Learning
In graph causal representation learning, researchers have explored various methods to improve the explanatory power and generalization performance of GNNs. By applying the principles of causal reasoning to graph-structured data, the researchers sought to address the challenges GNNs face when dealing with complex systems such as social networks, molecular maps, and syntax trees of program code. DIR Wu et al. (2022) is proposed to reveal the intrinsic interpretability of GNNs by discovering invariant reasons, which involves splitting input graphs into causal and non-causal fruit graphs and training the two classifiers through invariant risk loss functions. GOOD Chen et al. (2022b) improves the cross-domain generalization of graphs by distinguishing invariant subgraphs from other parts of graphs that are susceptible to domain transfer. CAL Sui et al. (2022) introduces de-confounding training to distinguish the key and secondary parts of the graph and eliminate the confounding effect of the secondary parts on model prediction. CFLP Zhao et al. (2022) points out that the causal relationship between graph structure and link presence is often ignored, and proposed to generate counterfactual links to enhance training data and reduce reliance on false associations. Zevcevic et al. Zečević et al. (2021) theoretically analyze the relationship between GNNs and structural causal models (SCMs) and design a new class of neuro-causal models. However, none of the work has been done to combine causal learning with the TKGR task.
@@ 3.1 Notations and Task Formulation
A TKG $\mathcal{G}$ can be formalized as a sequence of knowledge graph slices $\left\{\mathcal{G}_{1},\mathcal{G}_{2},\ldots.,\mathcal{G}_{T}\right\}$, where $\mathcal{G}_{t}=\left\{\left(e_{s},r,e_{o},t\right)\in\mathcal{G}\right\}$ denotes a knowledge graph slice that consists of facts that occurred at the timestamp $t$ range from $t_{0}$ to $t_{n}$. Here, $e_{s}$ and $e_{o}$ represent the subject and object entities, respectively, and $r$ denotes the predicate as a relation type. Besides, $\mathbf{e}_{s}$, $\mathbf{r}$, $\mathbf{e}_{o}$ written in bold represent their embeddings. The objective of TKGR task is to predict either the subject in a give query (?$,r,e_{o},t\right)$ or the object in a given query $\left(e_{s},r,?,t\right)$ with $t>t_{n}$.
3.2 A Causal Perspective on the GNN-Based TKGR Task
3.2.1 GNN-based TKGR Paradigm
Inspired by previous GNN-based modeling in a casual look Didelez and Pigeot (2001); Sui et al. (2022), we abstract the GNN-based TKGR process through a structural causal figure, as shown in Fig. 1, encompassing five distinct variables. The connectivity from one variable to another epitomizes the causal relationship, delineated as the cause $\rightarrow$ effect. The variables are described as follows:
Figure 1: The GNN-based structural causal graph for the TKGR task.
  • Graph data $\mathcal{G}_{t}$: The knowledge graph at each timestamp $t$, manifests as a directed multi-relationship figure.
@@ 3.2.1 GNN-based TKGR Paradigm
  • Prediction $Y$: Denoted as TKGR as the link prediction, this aspect transitions through the decoder, rendering the ultimate reasoning based on the preceding representation.
The causal embedding encapsulates the causal features $C$, authentically mirroring the implicit knowledge inherent in the knowledge graph $\mathcal{G}_{t}$. Conversely, $N$ symbolizes the confounding features, which might be spawned by data biases, data noise, or superficial patterns within graph-based learning methodologies. These confounding features forge a backdoor pathway between $C$ and $Y$, fostering spurious correlations that don’t contribute to accurate reasoning. Functionally, the structural operation denoted by $C$rightarrow $R$leftarrow $N$ portrays a GNN, wherein both the causal features $C$ and the confounding features $N$, as discerned by the target entity from the graph data, exert a direct impact on the output $R$ of the GNN. Subsequently, the output $R$ of GNN directly sways the model inference outcome, illustrated as $R$rightarrow $Y$.
In the graph-based TKGR paradigm, causal and confounding features are not decoupled for each entity or relationship embedding. Using causal graphs, we aspire to explicitly separate causal embeddings and confounding embeddings from entity or relational representations, and aim to mitigate the effects of confounding features by performing causal interventions. This endeavor not only clarifies the inference process but also endeavors to refine the accuracy and reliability of the GNN-based TKGR mechanism.
3.3 Causal Intervention Strategies
Beyond fostering a novel comprehension of GNN-based TKGR, causal theory avails analytical instruments predicated on causal figures, such as causal intervention. Causal intervention facilitates a profound examination of the factors precipitating inference outcomes. As delineated by Fig. 1., confounding feature $N$ and causal feature $C$ can be discerned from the knowledge graph $\mathcal{G}_{t}$. These features are contemplated in the representation $R$ of entities and relations, thereby establishing a backdoor pathway represented as $N$leftarrow $\mathcal{G}_{t}$rightarrow $C$rightarrow $R$rightarrow $Y$, with $N$ serving as the quick bridge between $C$ and $Y$.
@@ 3.3 Causal Intervention Strategies
$P(Y\mid do(C))=\sum_{n\in N}P(Y\mid C,n)P(n).$ (1)
The equation above illustrates that to gauge the causal influence of $C$ on $Y$, it’s requisite to take into account the inference outcomes of both causal and confounding features. This can be perceived as re-coupling the disentanglement feature embeddings, utilizing them for deductive reasoning at future timestamps. However, $C$ and $N$ are usually unobservable, and it is difficult to obtain them directly at the data level, which makes the calculation of the Eq. (1) very challenging. In the next section, we discuss ways to overcome this problem.
4 The Proposed CEGRL-TKGR Framework
4.1 The Overall Architecture of TKGR-GPRSCL
We detail the CEGRL-TKGR framework for learning representations of entities and relationships based on causal features and confounding features. CEGRL-TKGR consists of three parts: (1) The representation learning part that learns the structure dependence in each $\mathcal{G}_{t}$; (2) The decoupling learning part that learns the entity and relation representations; (3) The decoder part that is designed based on the time interval. The overall architecture of the framework is shown in Fig. 2.
Figure 2: The overall architecture of our proposed CEGRL-TKGR framework.
@@ 4.2 Entity and Relation Evolution Representation
Within each $\mathcal{G}_{t}$, representation learning of entities and relationships involves the aggregation of multiple relationships, as well as information from multiple hop neighbors under a single timestamp. Between adjacent $\mathcal{G}_{t}$, we expect to accurately capture the order dependencies inherent in the subgraph with different timestamps. Drawing inspiration from the RE-GCN model Li et al. (2021), we employ the $\omega$-layer RGCN, which hinges on structure modeling and a recurrent mechanism to progressively update the representations of entities and relations. This approach allows for a more nuanced understanding and modeling of the dynamic interactions within the graph over time.
$\mathbf{e}^{l+1}_{o,t}$ $=\operatorname{RReLu}\Bigg{(}\sum_{(e_{s},r,e_{o})\in\mathcal{G}_{t}}\frac{1}{d_{e_{o}}}\mathbf{W}^{l}_{1}\left(\Phi\left(\mathbf{e}^{l}_{s,t},\mathbf{r}_{t}\right)\right)$
$\qquad+\mathbf{W}^{l}_{2}\mathbf{e}^{l}_{o,t}\Bigg{)},$ (2)
$\mathbf{E}_{t}$ $=\operatorname{GRU}\left(\mathbf{E}_{t-1},\mathbf{E}^{\prime}_{t}\right).$ (3)
In the Eq. (2), we describe how the embedding $\mathbf{e}^{l+1}_{o,t}$ of entity $e_{o}$ at time step $t$ and layer $l+1$ is computed. We integrate the information of all entities and relations connected to entity $e_{o}$ in the knowledge graph $\mathcal{G}_{t}$. $\mathbf{W}^{l}_{1},\mathbf{W}^{l}_{2}$ is learnable weights and $\Phi$ has the option of addition or one-dimensional convolution. In the Eq. (3), we showcase how the entity embedding matrix $\mathbf{E}_{t}$ is updated via the GRU. Specifically, we take the entity embedding matrix $\mathbf{E}_{t-1}$ at the previous time step $t-1$ and the aggregated entity embedding matrix $\mathbf{E}^{\prime}_{t}$ as inputs to obtain the entity embedding matrix $\mathbf{E}_{t}$ at the current time step $t$.
@@
$\mathbf{R}_{t}$ $=\operatorname{GRU}\left(\mathbf{R}_{t-1},\mathbf{R}^{\prime}_{t}\right),$ (5)
where $\mathbf{r}^{\prime}_{t}$ is an aggregation of all entities connected to relation $r$ via a mean pooling operation, and $\mathbf{R}^{\prime}_{t}$ is obtained by concatenating this result with the embeddings of all relations. Eventually, we update the relation embedding matrix $\mathbf{R}_{t}$ using a GRU.